# Safe Multi-Robotic Arm Interaction via 3D Convex Shapes

### v2

@@ Abstract
Inter-robot collisions pose a significant safety risk when multiple robotic arms operate in close proximity. We present an online collision avoidance methodology leveraging 3D convex shape-based High-Order Control Barrier Functions (HOCBFs) to address this issue. While prior works focused on using Control Barrier Functions (CBFs) for human–robotic arm and single-arm collision avoidance, we explore the problem of collision avoidance between multiple robotic arms operating in a shared space. In our methodology, we utilize the proposed HOCBFs as centralized and decentralized safety filters. These safety filters are compatible with many nominal controllers and ensure safety without significantly restricting the robots’ workspace. A key challenge in implementing these filters is the computational overhead caused by the large number of safety constraints and the computation of a Hessian matrix per constraint. We address this challenge by employing numerical differentiation methods to approximate computationally intensive terms. The effectiveness of our method is demonstrated through extensive simulation studies and real-world experiments with Franka Research 3 robotic arms. The project video is available at this link.
@@ 1 Introduction
The deployment of robots in unstructured dynamic environments necessitates effective collision prevention mechanisms. Various approaches have been proposed for this purpose, including Model Predictive Control (MPC)-based safety strategies [1, 2] and CBF-based methods [3, 4, 5, 6, 7, 8]. CBF-based methods have received attention because of their ability to enforce forward invariance of safety sets while maintaining computational efficiency in many scenarios. Leveraging these properties, recent research has extensively explored their application in inter-robot collision avoidance, particularly for safe navigation in mobile robot systems [4, 5, 6, 7]. However, despite the importance of multi-robotic arm interaction in specific applications [9, 10], CBF-based safety certificates for ensuring safe multi-robotic arm interaction remain largely unexplored. This work addresses this gap by investigating the application of CBFs that enhance safety during multi-robotic arm interactions and collaborative tasks.
CBFs are often used as constraints in Quadratic Programming (QP) to enforce safety [11]. The QP aims to keep the control input as close to the nominal controller as possible while obeying the safety constraints. To generalize this framework for systems whose relative degree with respect to the constraint function is higher than one, HOCBFs were introduced [12, 13]. Our methodology is built on this concept by employing HOCBFs for collision avoidance among 3D convex shapes, as introduced in [3]. We extend the use of HOCBFs from single robotic arm collision avoidance to inter-robotic arm collision avoidance by utilizing them in pairwise robotic arm dynamics.
The key challenge in utilizing 3D Convex Shape-based HOCBFs in multi-robotic arms, compared to a single arm, is the increased computational overhead. This arises from calculating more constraints at every sampling time. Additionally, in multi-robot settings, all the 3D convex shapes used in the HOCBFs are dynamic. This further intensifies the computational burden. We address this issue by approximating computationally intensive terms via numerical derivative filters. Our contributions in this paper are summarized as follows:
@@ 2 Related works
CBFs in inter-robot collision avoidance: The use of CBFs to prevent collisions in multi-robot systems is an active research area, with various methodologies proposed for different scenarios [4, 5, 6, 7, 8]. For instance, centralized and decentralized CBF-based safety filters for planar mobile robots were introduced in [4]. A CBF-based methodology was developed for the safe control of drone swarms in [6, 14]. In addition, a distributed controller for the safe navigation of mobile robots was proposed in [8]. Similarly, HOCBFs were employed to generate safety-assured trajectories for multiple legged robots in [5]. However, extending these approaches to the multi-robotic arm setting is not straightforward. Although some work [15, 16] applied CBFs to ensure safe human-multiple robot arm interaction, they did not address possible collisions between robot bodies. This is the first work applying CBFs for collision avoidance among multiple robotic arms.
Inter-robotic arm collision avoidance: Collision handling between robotic arms can be categorized into online [17, 18, 19] and offline [20, 21] methods. Offline methods predefine commands and trajectories, while online methods address collisions in real time. Various online strategies have been explored. For example, collision maps were utilized to detect potential collisions and schedule command execution times while preserving the path in [18]. Nonlinear MPC was employed for online trajectory planning, incorporating collision avoidance into the optimization constraints in [19]. Additionally, a motion planning method based on Soft Actor-Critic was proposed in [17] to prevent collisions in dual-robotic arms.
@@ 4.1 High order control barrier functions
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times q}$ are Lipschitz $C^{1}$ functions with the state vector $\mathbf{x}\in\mathbb{R}^{n}$ and input vector $\mathbf{u}\in\mathcal{U}\subset\mathbb{R}^{q}$, where $\mathcal{U}$ denotes a closed set of possible control input values (taking into account physical constraints). For a given $C^{r}$ function $h(\mathbf{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$ and $r$ class $\mathcal{K}$ functions $\Gamma_{k}(\cdot)$, $k\in\{1,\ldots,r\}$, a sequence of functions and their superlevel sets are defined as follows:
@@
$\psi_{k}(\mathbf{x})$ $=\left(\frac{d}{dt}+\Gamma_{k}\right)\psi_{k-1}(\mathbf{x}),\quad k\in\{1,\ldots,r\},$
@@
Definition 1 (Time-invariant HOCBFs [13]).
@@
Theorem 1 (Theorem 4 in [13] for time-invariant HOCBFs).
@@ 4.2 Convex shapes for defining safety in robotic arms
Definition 2 (Scaling functions [22]).
@@
$\mathbf{p}\in\partial A(\boldsymbol{\theta})\quad\Longleftrightarrow\quad\mathcal{F}_{A}(\mathbf{p},\boldsymbol{\theta})=1.$ (6)
Let $A$ and $B$ be convex sets in $\mathbb{R}^{n_{p}}$ with $A\cap B=\emptyset$, and let their corresponding scaling functions be $\mathcal{F}_{A}(\mathbf{p},\boldsymbol{\theta})$ and $\mathcal{F}_{B}(\mathbf{p},\boldsymbol{\theta})$, respectively. In [3], determining the minimum level set of $\mathcal{F}_{A}(\mathbf{p},\boldsymbol{\theta})$ that intersects $B$ is formulated as the following optimization problem:
@@
Although there are alternative formulations, we choose the approach in [3] because it facilitates the use of torque control in robotic arms. Let $\mathbf{p}^{\star}(\boldsymbol{\theta})$ denote the optimal solution to this problem. The following theorem provides sufficient conditions for the differentiability order of $\alpha^{\star}(\boldsymbol{\theta})$:
Theorem 2 (Theorem 2 in [3]).
Let $\mathcal{F}_{A}$ and $\mathcal{F}_{B}$ be scaling functions corresponding to the sets $A$ and $B$, respectively, with $A\cap B=\emptyset$. Assume that one of $\mathcal{F}_{A}$ and $\mathcal{F}_{B}$ has a positive definite Hessian in $\mathbf{p}$. If the scaling functions $\mathcal{F}_{A}$ and $\mathcal{F}_{B}$ are $C^{k+1}$ (with $k\geq 1$) in $\mathbf{p}$ and $\boldsymbol{\theta}$, then $\alpha^{\star}$ is $C^{k}$ in $\boldsymbol{\theta}$.
@@ 5.1 Construction of HOCBFs for multi-robotic arm setting
Consider an environment with several robotic arms and each robotic arm in the environment is indexed by the set $\mathcal{I}\subset\mathbb{N}$. Dynamics of robotic arm $i\in\mathcal{I}$ are written as [23]:
$M_{i}(\mathbf{q}_{i})\ddot{\mathbf{q}}_{i}+\underbrace{C_{i}(\mathbf{q}_{i},\dot{\mathbf{q}}_{i})\dot{\mathbf{q}}_{i}+F_{i}(\dot{\mathbf{q}}_{i})+g_{i}(\mathbf{q}_{i})}_{\sigma_{i}(\mathbf{q}_{i},\dot{\mathbf{q}}_{i})}=\mathbf{u}_{i}$ (8)
@@
where $\mathbf{q}_{i}\in\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ and $\mathbf{u}_{i}\in\mathcal{U}_{i}\subset\mathbb{R}^{n_{i}}$ are the joint vector and actuator torque vector for robot $i$, respectively. The sets $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ represent the closed sets of the possible joint configurations and control inputs for robot $i$. $M_{i}(\mathbf{q}_{i})\in\mathbb{R}^{n_{i}\times n_{i}}$ is the symmetric positive definite mass matrix, $C_{i}(\mathbf{q}_{i},\dot{\mathbf{q}}_{i})\in\mathbb{R}^{n_{i}\times n_{i}}$ consists of centrifugal and Coriolis terms, $F_{i}(\dot{\mathbf{q}}_{i})\in\mathbb{R}^{n_{i}}$ is the friction vector, and $g_{i}(\mathbf{q}_{i})\in\mathbb{R}^{n_{i}}$ is the gravity vector for robot $i$. By utilizing (8), the pairwise robot dynamics for two different robot $i$ and $j$ ($i<j$ to impose a unique ordering convention on pairs) can be expressed in a control affine system form as follows:
$\frac{d}{dt}\underbrace{\begin{bmatrix}\mathbf{q}_{i}\\ \dot{\mathbf{q}}_{i}\\ \mathbf{q}_{j}\\ \dot{\mathbf{q}}_{j}\end{bmatrix}}_{\mathbf{x}_{ij}}=$ $\underbrace{\begin{bmatrix}\dot{\mathbf{q}}_{i}\\ M^{-1}_{i}(\mathbf{q}_{i})\sigma_{i}(\mathbf{q}_{i},\dot{\mathbf{q}_{i}})\\ \dot{\mathbf{q}}_{j}\\ M^{-1}_{j}(\mathbf{q}_{j})\sigma_{j}(\mathbf{q}_{j},\dot{\mathbf{q}_{j}})\end{bmatrix}}_{f(\mathbf{x}_{ij})}+\underbrace{\begin{bmatrix}0&0\\ M^{-1}_{i}(\mathbf{q}_{i})&0\\ 0&0\\ 0&M^{-1}_{j}(\mathbf{q}_{j})\end{bmatrix}}_{g(\mathbf{x}_{ij})}\underbrace{\begin{bmatrix}\mathbf{u}_{i}\\ \mathbf{u}_{j}\end{bmatrix}}_{\mathbf{u}_{ij}}.$ (9)
@@
To construct a scaling function for a robot’s link or end effector, we follow the construction recipe described in [3]. For a given $^{\scriptscriptstyle W}\mathbf{p}\in\mathbb{R}^{3}$ in the world frame, the point is first transformed into body frame coordinates as $^{\scriptscriptstyle B}\mathbf{p}={}^{\scriptscriptstyle B}\mathbf{R}_{\scriptscriptstyle W}({}^{\scriptscriptstyle W}\mathbf{p}-{}^{\scriptscriptstyle W}\mathbf{o}_{\scriptscriptstyle B})$ where $^{\scriptscriptstyle B}\mathbf{R}_{\scriptscriptstyle W}$ denotes the orientation of the world frame expressed in the body frame, and $^{\scriptscriptstyle W}\mathbf{o}_{\scriptscriptstyle B}$ represents the origin of the body frame in the world frame.
@@
An analytical solution for (7) is available when both scaling functions are ellipsoids [24]. In contrast, computing the HOCBF coefficients for other 3D convex shapes requires numerically solving (7) at every iteration. This process is computationally costly for real-time robot control. Therefore, we use ellipsoids to represent the robot links and end effector. The scaling functions used in the remainder of this paper are of the following form:
@@
$\psi_{0}^{i_{k}j_{l}}(\mathbf{x}_{ij})=h_{i_{k}j_{l}}(\mathbf{x}_{ij})=\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})-\alpha_{0},\>\>\alpha_{0}>1$ (11)
$\psi_{w}^{i_{k}j_{l}}(\mathbf{x}_{ij})=\left(\frac{d}{dt}+\Gamma_{w}\right)\psi_{w-1}^{i_{k}j_{l}}(\mathbf{x}_{ij}),\quad w\in\{1,2\}.$
@@
$\mathbf{T}_{ij}^{kl}=\begin{bmatrix}\mathbf{I}_{3\times 3}&0&0&0\\ 0&0.5Q(\boldsymbol{\xi}_{i_{k}})&0&0\\ 0&0&\mathbf{I}_{3\times 3}&0\\ 0&0&0&0.5Q(\boldsymbol{\xi}_{j_{l}})\end{bmatrix}$, (14)
@@
with the matrix $Q(\boldsymbol{\xi})$ defined as
$Q(\boldsymbol{\xi})=\begin{bmatrix}\boldsymbol{\xi}_{w}&\boldsymbol{\xi}_{z}&-\boldsymbol{\xi}_{y}\\ -\boldsymbol{\xi}_{z}&\boldsymbol{\xi}_{w}&\boldsymbol{\xi}_{x}\\ \boldsymbol{\xi}_{y}&-\boldsymbol{\xi}_{x}&\boldsymbol{\xi}_{w}\\ -\boldsymbol{\xi}_{x}&-\boldsymbol{\xi}_{y}&-\boldsymbol{\xi}_{z}\end{bmatrix}$, (15)
@@
$+\left(\nabla_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\right)^{T}\left(\dot{\mathbf{T}}_{ij}^{kl}\boldsymbol{\Omega}_{ij}^{kl}\dot{\mathbf{q}}_{ij}+\mathbf{T}_{ij}^{kl}\dot{\boldsymbol{\Omega}}_{ij}^{kl}\dot{\mathbf{q}}_{ij}+\mathbf{T}_{ij}^{kl}\boldsymbol{\Omega}_{ij}^{kl}\boldsymbol{\Upsilon}_{ij}\boldsymbol{\sigma}_{ij}\right),$

### v3

@@ Abstract
Inter-robot collisions pose a significant safety risk when multiple robotic arms operate in close proximity. We present an online collision avoidance methodology leveraging High-Order Control Barrier Functions (HOCBFs) constructed for safe interactions among 3D convex shapes to address this issue. While prior works focused on using Control Barrier Functions (CBFs) for human–robotic arm and single-arm collision avoidance, we explore the problem of collision avoidance between multiple robotic arms operating in a shared space. In our methodology, we utilize the proposed HOCBFs as centralized and decentralized safety filters. These safety filters are compatible with many nominal controllers and ensure safety without significantly restricting the robots’ workspace. A key challenge in implementing these filters is the computational overhead caused by the large number of safety constraints and the computation of a Hessian matrix per constraint. We address this challenge by employing numerical differentiation methods to approximate computationally intensive terms. The effectiveness of our method is demonstrated through extensive simulation studies and real-world experiments with Franka Research 3 robotic arms. The project video is available at this link. © 2025. This manuscript version is made available under the CC-BY-NC-ND 4.0 license https://creativecommons.org/licenses/by-nc-nd/4.0/.
The formal publication is available at DOI: https://doi.org/10.1016/j.robot.2025.105263.
@@ 1 Introduction
The deployment of robots in unstructured dynamic environments necessitates effective collision prevention mechanisms. Various approaches have been proposed for this purpose, including Model Predictive Control (MPC)-based safety strategies sathya2018embedded , zhang2020optimization and CBF-based methods wei2024collision , wang2017safety , kim2024safety , GONCALVES2024104601 , cavorsi2023multi , mestres2024distributed . CBF-based methods have received attention because of their ability to enforce forward invariance of safety sets while maintaining computational efficiency in many scenarios. Leveraging these properties, recent research has extensively explored their application in inter-robot collision avoidance, particularly for safe navigation in mobile robot systems wang2017safety , kim2024safety , GONCALVES2024104601 , cavorsi2023multi . However, despite the importance of multi-robotic arm interaction in specific applications smith2012dual , yan2021decentralized , CBF-based safety certificates for ensuring safe multi-robotic arm interaction remain largely unexplored. This work addresses this gap by investigating the application of CBFs that enhance safety during multi-robotic arm interactions and collaborative tasks.
CBFs are often used as constraints in Quadratic Programming (QP) to enforce safety ames2019control . The QP aims to keep the control input as close to the nominal controller as possible while obeying the safety constraints. To generalize this framework for systems whose relative degree with respect to the constraint function is higher than one, HOCBFs were introduced tan2021high , xiao2021high . Our methodology is built on this concept by employing HOCBFs for collision avoidance among 3D convex shapes, as introduced in wei2024collision . We extend the use of HOCBFs from single robotic arm collision avoidance to inter-robotic arm collision avoidance by utilizing them in pairwise robotic arm dynamics.
The key challenge in utilizing HOCBFs constructed for safe interactions among 3D convex shapes in multi-robotic arms, compared to a single arm, is the increased computational overhead. This arises from calculating more constraints at every sampling time. Additionally, in multi-robot settings, all the 3D convex shapes used in the HOCBFs are dynamic. This further intensifies the computational burden. We address this issue by approximating computationally intensive terms via numerical derivative filters. Our contributions in this paper are summarized as follows:
@@ 2 Related works
CBFs in inter-robot collision avoidance: The use of CBFs to prevent collisions in multi-robot systems is an active research area, with various methodologies proposed for different scenarios wang2017safety , kim2024safety , GONCALVES2024104601 , cavorsi2023multi , mestres2024distributed . For instance, centralized and decentralized CBF-based safety filters for planar mobile robots were introduced in wang2017safety . A CBF-based methodology was developed for the safe control of drone swarms in GONCALVES2024104601 , GKT24 . In addition, a distributed controller for the safe navigation of mobile robots was proposed in mestres2024distributed . Similarly, HOCBFs were employed to generate safety-assured trajectories for multiple legged robots in kim2024safety . However, extending these approaches to the multi-robotic arm setting is not straightforward. Although some work shi2024safe , shi2025distributed applied CBFs to ensure safe human-multiple robot arm interaction, they did not address possible collisions between robot bodies. This is the first work applying CBFs for collision avoidance among multiple robotic arms.
Inter-robotic arm collision avoidance: Collision handling between robotic arms can be categorized into online wong2021motion , afaghani2015line , gafur2022dynamic and offline chang1994collision , chen2024multiobjective methods. Offline methods predefine commands and trajectories, while online methods address collisions in real time. Various online strategies have been explored. For example, collision maps were utilized to detect potential collisions and schedule command execution times while preserving the path in afaghani2015line . Nonlinear MPC was employed for online trajectory planning, incorporating collision avoidance into the optimization constraints in gafur2022dynamic . Additionally, a motion planning method based on Soft Actor-Critic was proposed in wong2021motion to prevent collisions in dual-robotic arms.
@@ 4.1 High order control barrier functions
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ and $g:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n\times q}$ are Lipschitz $\mathcal{C}^{1}$ functions with the state vector $\mathbf{x}\in\mathbb{R}^{n}$ and input vector $\mathbf{u}\in\mathcal{U}\subset\mathbb{R}^{q}$, where $\mathcal{U}$ denotes a closed set of possible control input values (taking into account physical constraints). For a given $\mathcal{C}^{r}$ function $h(\mathbf{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$ and $r$ class $\mathcal{K}$ functions $\Gamma_{k}(\cdot)$, $k\in\{1,\ldots,r\}$, a sequence of functions and their superlevel sets are defined as follows:
@@
$\psi_{k}(\mathbf{x})$ $=\frac{d}{dt}\psi_{k-1}(\mathbf{x})+\Gamma_{k}\left(\psi_{k-1}(\mathbf{x})\right),\quad k\in\{1,\ldots,r\},$
@@
Definition 1 (Time-invariant HOCBFs xiao2021high ).
@@
Theorem 1 (Theorem 4 in xiao2021high for time-invariant HOCBFs).
@@ 4.2 Convex shapes for defining safety in robotic arms
Definition 2 (Scaling functions wei2024diffocclusion ).
@@
$\mathbf{p}\in\partial A(\boldsymbol{\theta})\quad\Longrightarrow\quad\mathcal{F}_{A}(\mathbf{p},\boldsymbol{\theta})=1.$ (6)
Let $A$ and $B$ be convex sets in $\mathbb{R}^{n_{p}}$ with $A\cap B=\emptyset$, and let their corresponding scaling functions be $\mathcal{F}_{A}(\mathbf{p},\boldsymbol{\theta})$ and $\mathcal{F}_{B}(\mathbf{p},\boldsymbol{\theta})$, respectively. In wei2024collision , determining the minimum level set of $\mathcal{F}_{A}(\mathbf{p},\boldsymbol{\theta})$ that intersects $B$ is formulated as the following optimization problem:
@@
Although there are alternative formulations, we choose the approach in wei2024collision because it facilitates the use of torque control in robotic arms. Let $\mathbf{p}^{\star}(\boldsymbol{\theta})$ denote the optimal solution to (7), i.e., the argument that solves (7). The following theorem provides sufficient conditions for the differentiability order of $\alpha^{\star}(\boldsymbol{\theta})$:
Theorem 2 (Theorem 2 in wei2024collision ).
Let $\mathcal{F}_{A}$ and $\mathcal{F}_{B}$ be scaling functions corresponding to the sets $A$ and $B$, respectively, with $A\cap B=\emptyset$. Assume that one of $\mathcal{F}_{A}$ and $\mathcal{F}_{B}$ has a positive definite Hessian in $\mathbf{p}$. If the scaling functions $\mathcal{F}_{A}$ and $\mathcal{F}_{B}$ are $\mathcal{C}^{k+1}$ (with $k\geq 1$) in $\mathbf{p}$ and $\boldsymbol{\theta}$, then $\alpha^{\star}$ is $\mathcal{C}^{k}$ in $\boldsymbol{\theta}$.
@@ 5.1 Construction of HOCBFs for multi-robotic arm setting
Consider an environment with several robotic arms and each robotic arm in the environment is indexed by the set $\mathcal{I}\subset\mathbb{N}$. Dynamics of robotic arm $i\in\mathcal{I}$ are written as siciliano2008robotics :
$M_{i}(\mathbf{q}_{i})\ddot{\mathbf{q}}_{i}+\underbrace{D_{i}(\mathbf{q}_{i},\dot{\mathbf{q}}_{i})\dot{\mathbf{q}}_{i}+F_{i}(\dot{\mathbf{q}}_{i})+g_{i}(\mathbf{q}_{i})}_{\sigma_{i}(\mathbf{q}_{i},\dot{\mathbf{q}}_{i})}=\mathbf{u}_{i}$ (8)
@@
where $\mathbf{q}_{i}\in\mathcal{X}_{i}\subset\mathbb{R}^{n_{i}}$ and $\mathbf{u}_{i}\in\mathcal{U}_{i}\subset\mathbb{R}^{n_{i}}$ are the joint vector and actuator torque vector for robot $i$, respectively. The sets $\mathcal{X}_{i}$ and $\mathcal{U}_{i}$ represent the closed sets of the possible joint configurations and control inputs for robot $i$. $M_{i}(\mathbf{q}_{i})\in\mathbb{R}^{n_{i}\times n_{i}}$ is the symmetric positive definite mass matrix, $D_{i}(\mathbf{q}_{i},\dot{\mathbf{q}}_{i})\in\mathbb{R}^{n_{i}\times n_{i}}$ consists of centrifugal and Coriolis terms, $F_{i}(\dot{\mathbf{q}}_{i})\in\mathbb{R}^{n_{i}}$ is the friction vector, and $g_{i}(\mathbf{q}_{i})\in\mathbb{R}^{n_{i}}$ is the gravity vector for robot $i$. By utilizing (8), the pairwise robot dynamics for two different robot $i$ and $j$ ($i<j$ to impose a unique ordering convention on pairs) can be expressed in a control affine system form as follows:
$\frac{d}{dt}\underbrace{\begin{bmatrix}\mathbf{q}_{i}\\ \dot{\mathbf{q}}_{i}\\ \mathbf{q}_{j}\\ \dot{\mathbf{q}}_{j}\end{bmatrix}}_{\mathbf{x}_{ij}}=$ $\underbrace{\begin{bmatrix}\dot{\mathbf{q}}_{i}\\ -M^{-1}_{i}(\mathbf{q}_{i})\sigma_{i}(\mathbf{q}_{i},\dot{\mathbf{q}_{i}})\\ \dot{\mathbf{q}}_{j}\\ -M^{-1}_{j}(\mathbf{q}_{j})\sigma_{j}(\mathbf{q}_{j},\dot{\mathbf{q}_{j}})\end{bmatrix}}_{f(\mathbf{x}_{ij})}+\underbrace{\begin{bmatrix}0&0\\ M^{-1}_{i}(\mathbf{q}_{i})&0\\ 0&0\\ 0&M^{-1}_{j}(\mathbf{q}_{j})\end{bmatrix}}_{g(\mathbf{x}_{ij})}\underbrace{\begin{bmatrix}\mathbf{u}_{i}\\ \mathbf{u}_{j}\end{bmatrix}}_{\mathbf{u}_{ij}}.$ (9)
@@
To construct a scaling function for a robot’s link or end effector, we follow the construction recipe described in wei2024collision . For a given $^{\scriptscriptstyle W}\mathbf{p}\in\mathbb{R}^{3}$ in the world frame, the point is first transformed into body frame coordinates as $^{\scriptscriptstyle B}\mathbf{p}={}^{\scriptscriptstyle B}\mathbf{R}_{\scriptscriptstyle W}({}^{\scriptscriptstyle W}\mathbf{p}-{}^{\scriptscriptstyle W}\mathbf{o}_{\scriptscriptstyle B})$ where $^{\scriptscriptstyle B}\mathbf{R}_{\scriptscriptstyle W}$ denotes the orientation of the world frame expressed in the body frame, and $^{\scriptscriptstyle W}\mathbf{o}_{\scriptscriptstyle B}$ represents the origin of the body frame in the world frame.
@@
An analytical solution for (7) is available when both scaling functions are ellipsoids rimon1997obstacle . In contrast, computing the HOCBF coefficients for other 3D convex shapes requires numerically solving (7) at every iteration. This process is computationally costly for real-time robot control. Therefore, we use ellipsoids to represent the robot links and end effector. The scaling functions used in the remainder of this paper are of the following form:
@@
$\psi_{0}^{i_{k}j_{l}}(\mathbf{x}_{ij})$ $=h_{i_{k}j_{l}}(\mathbf{x}_{ij})=\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})-\alpha_{0},\>\>\alpha_{0}>1,$ (11)
$\psi_{w}^{i_{k}j_{l}}(\mathbf{x}_{ij})$ $=\frac{d}{dt}\psi_{w-1}^{i_{k}j_{l}}(\mathbf{x}_{ij})+\Gamma_{w}\left(\psi_{w-1}^{i_{k}j_{l}}(\mathbf{x}_{ij})\right),\quad w\in\{1,2\}.$
By the definition of $\alpha^{\star}(\boldsymbol{\theta})$ in (7), the collision bodies of the pair $(i_{k},j_{l})$ do not intersect if $\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})$ is greater than $1$. Therefore, setting the parameter $\alpha_{0}>1$ in (11) ensures that the pair $(i_{k},j_{l})$ of robot links or end-effectors remain collision-free within the safe set defined by the proposed HOCBF. Specifically, the parameter $\alpha_{0}$ defines the safety margin between these bodies by shifting the boundary of the feasible region. Larger values of $\alpha_{0}$ correspond to more conservative safety distances, while smaller values allow robots to operate closer to each other.
@@
$\mathbf{T}_{ij}^{kl}=\begin{bmatrix}\mathbf{I}_{3\times 3}&0&0&0\\ 0&0.5\Pi(\boldsymbol{\xi}_{i_{k}})&0&0\\ 0&0&\mathbf{I}_{3\times 3}&0\\ 0&0&0&0.5\Pi(\boldsymbol{\xi}_{j_{l}})\end{bmatrix}$, (14)
@@
with the matrix $\Pi(\boldsymbol{\xi})$ defined as
$\Pi(\boldsymbol{\xi})=\begin{bmatrix}\boldsymbol{\xi}_{w}&\boldsymbol{\xi}_{z}&-\boldsymbol{\xi}_{y}\\ -\boldsymbol{\xi}_{z}&\boldsymbol{\xi}_{w}&\boldsymbol{\xi}_{x}\\ \boldsymbol{\xi}_{y}&-\boldsymbol{\xi}_{x}&\boldsymbol{\xi}_{w}\\ -\boldsymbol{\xi}_{x}&-\boldsymbol{\xi}_{y}&-\boldsymbol{\xi}_{z}\end{bmatrix}$, (15)
@@
$+\left(\nabla_{\boldsymbol{\theta}}\alpha^{\star}(\boldsymbol{\theta}_{i_{k}j_{l}})\right)^{T}\left(\dot{\mathbf{T}}_{ij}^{kl}\boldsymbol{\Omega}_{ij}^{kl}\dot{\mathbf{q}}_{ij}+\mathbf{T}_{ij}^{kl}\dot{\boldsymbol{\Omega}}_{ij}^{kl}\dot{\mathbf{q}}_{ij}-\mathbf{T}_{ij}^{kl}\boldsymbol{\Omega}_{ij}^{kl}\boldsymbol{\Upsilon}_{ij}\boldsymbol{\sigma}_{ij}\right),$